# Algebraic and geometric properties of homeomorphism groups of ordinals

### v1

@@ 1 Introduction
An ordinal has a natural topology given by the order topology. Motivated by recent results in the study of homeomorphism groups of 2-manifolds, we investigate several algebraic and geometric properties of the homeomorphisms group of ordinals. We provide a discussion of this motivation from geometric topology after introducing our main results.
For an introduction to ordinals and the terminology that follows, see Section 2. Our results focus on successor ordinals, or equivalently, the compact ordinals. Up to homeomorphism, a successor ordinal has the form $\omega^{\alpha}\cdot d+1$, where $\alpha$ is an ordinal (called the limit capacity), $\omega$ is the first infinite ordinal, and $d\in\omega$ (called the coefficient). We study the successor ordinals for which the limit capacity is a successor ordinal. Given an ordinal $\alpha$ and $d\in\omega$, let $\operatorname{H}_{\alpha,d}=\operatorname{Homeo}(\omega^{\alpha+1}\cdot d+1)$, i.e., the group consisting of homeomorphisms $\omega^{\alpha+1}\cdot d+1\to\omega^{\alpha+1}\cdot d+1$.
Our first slate of results is for the case when the coefficient is one. In this case, $\operatorname{H}_{\alpha,1}$ is isomorphic to $\operatorname{Homeo}(\omega^{\alpha+1})$, and so we state our results for $\operatorname{Homeo}(\omega^{\alpha+1})$, as this is the view our proofs will take. This class of groups can be thought of as a generalization of $\operatorname{Sym}(\mathbb{N})$, the symmetric group on the natural numbers, as $\operatorname{Sym}(\mathbb{N})$ is isomorphic to $\operatorname{Homeo}(\omega)$.
An element $g$ in a group $G$ is a normal generator of $G$ if every element of $G$ can be expressed as product of conjugates of $g$ and $g^{-1}$; it is a uniform normal generator of $G$ if there exists $k\in\mathbb{N}$ [^1] Throughout, we use the convention that $0\notin\mathbb{N}$, i.e., $\mathbb{N}=\omega\smallsetminus\{0\}$. such that every element of $G$ can be expressed a product of at most $k$ conjugates of $g$ and $g^{-1}$, and the minimum such $k$ is called the $g$-width of $G$.
Every element of an ordinal is itself an ordinal. From this, we say an element of $\omega^{\alpha+1}$ is maximal rank if it is of the form $\omega^{\alpha}\cdot k$ for some $k\in\mathbb{N}$.
@@
    $h$ normally generates $\operatorname{Homeo}(\omega^{\alpha+1})$.
@@
Moreover, if one—and hence all—of the above conditions are satisfied, then the $h$-width of $\operatorname{Homeo}(\omega^{\alpha+1})$ is at most twelve.
The proof of Theorem 3.12 relies on a uniform fragmentation result—a generalization of a lemma of Galvin [14] from permutation groups—and a technique from Anderson’s proof [1] of the algebraic simplicity of several transformation groups, including the homeomorphism group of the Cantor set and the homeomorphism group of the 2-sphere. As an immediate consequence, $\operatorname{Homeo}(\omega^{\alpha+1})$ has a maximal proper normal subgroup.
@@
Let $\alpha$ be an ordinal. The subgroup $\mathrm{Fin}(\omega^{\alpha+1})$ of $\operatorname{Homeo}(\omega^{\alpha+1})$ consisting of homeomorphisms that induce a finite permutation on the set of maximal rank elements in $\omega^{\alpha+1}$ is the maximal proper normal subgroup of $\operatorname{Homeo}(\omega^{\alpha+1})$, that is, it contains every proper normal subgroup. ∎
As $\operatorname{Homeo}(\omega)$ is isomorphic to $\operatorname{Sym}(\mathbb{N})$, 1.1 recovers the classical theorem of Schreier–Ulam [27] that the subgroup of finite permutations in $\operatorname{Sym}(\mathbb{N})$ is its maximal normal subgroup. The proof presented here is arguably simpler than the standard proof given in classical texts such as in [28], which relies on cycle decompositions. Similarly, 3.12 can be viewed as a generalization of a theorem of Bertram [6] showing that all infinite permutations in $\operatorname{Sym}(\mathbb{N})$ are uniform normal generators. Unlike Bertram’s proof, our proof does not rely on cycle decompositions. Bertram proves that any permutation in $\operatorname{Sym}(\mathbb{N})$ can be expressed as a product of four conjugates of any infinite permutation, raising the question of what the optimal constant in 3.12 is.
A group $G$ is perfect if it is equal to its commutator subgroup $[G,G]$, i.e., the subgroup generated by the set $\{[x,y]:x,y\in G\}$, where $[x,y]=xyx^{-1}y^{-1}$. It is uniformly perfect if there exists some $k\in\mathbb{N}$ such that every element of $G$ can be expressed as a product of $k$ commutators; the minimum such $k$ is called the commutator width of $G$.
It is not difficult to find a commutator in $\operatorname{Homeo}(\omega^{\alpha+1})$ that induces an infinite permutation on the set of maximal rank elements of $\omega^{\alpha+1}$; 3.12 then implies that $\operatorname{Homeo}(\omega^{\alpha+1})$ is a uniformly perfect group of commutator width at most twelve. A direct proof, which is simpler than the proof of 3.12, can be given of this fact that provides a better bound on the commutator width of $G$. This is our second main result.
@@
A theorem of Ore [24] says that every element of $\operatorname{Sym}(\mathbb{N})$ can be expressed as a single commutator. It is natural to ask whether the same is true for $\operatorname{Homeo}(\omega^{\alpha+1})$.
Continuing in analogy with $\operatorname{Sym}(\mathbb{N})$, we turn to a geometric property. In [5], Bergman showed that any left-invariant metric on $\operatorname{Sym}(\mathbb{N})$ has bounded diameter; in the literature, this property is referred to as strong boundedness or the Bergman property. There are several equivalent definitions of strongly bounded, one of relevance here is that every group action on a metric space has bounded orbits. As a consequence, any homomorphism from a strongly bounded group to a countable group has finite image, a fact we use below.
Adapting a technique appearing several places in the literature (see [21, Construction 2.3]), we establish a property stronger than strong boundedness for $\operatorname{Homeo}(\omega^{\alpha+1})$. A group $G$ is strongly distorted if there exists $m\in\mathbb{N}$ and $\{w_{n}\}_{n\in\mathbb{N}}\subset\mathbb{N}$ such that for any sequence $\{g_{n}\}_{n\in\mathbb{N}}$ in $G$ there exists $S\subset G$ of cardinality at most $m$ satisfying $g_{n}\in S^{w_{n}}$, where $S^{k}=\{s_{1}s_{2}\cdots s_{k}:s_{1},\ldots,s_{k}\in S\}$. The notion of strong distortion was introduced by Calegari–Freedman in [9], where they established strong distortion for the homeomorphism groups of spheres; in the appendix of the same paper, Cornulier showed that if $G$ is strongly distorted, then $G$ is strongly bounded.
@@
We finish our discussion of the degree one case by observing that we cannot drop the hypothesis that the limit capacity is a successor. It is a consequence of [12, Theorem 3.9] that every countable group is a quotient of $\operatorname{Homeo}(\omega_{1})$, where $\omega_{1}$ is the first uncountable ordinal. As a consequence, $\operatorname{Homeo}(\omega_{1})$ can be neither uniformly perfect, strongly bounded, nor strongly distorted.
We now turn to the case of coefficient, or equivalently Cantor–Bendixson degree, greater than one, investigating the algebraic structure of $\operatorname{H}_{\alpha,d}$ and establishing the failure of the above results when $d>1$. There are several normal subgroups of $\operatorname{H}_{\alpha,d}$ that are essential to our results, which we must introduce. The first is $\operatorname{F}_{\alpha,d}$, the subgroup consisting of the homeomorphisms of $\omega^{\alpha+1}\cdot d+1$ that induce a finite permutation on the subset $N_{\alpha,d}:=\{\omega^{\alpha+1}\cdot k+\omega^{\alpha}\cdot\ell:k\in d,\ell\in\mathbb{N}\}$, i.e., the elements of next-to-maximal rank. The second is obtained by equipping $\operatorname{H}_{\alpha,d}$ with the compact-open topology and taking the closure of $\operatorname{F}_{\alpha,d}$, which we denote $\overline{\operatorname{F}}_{\alpha,d}$.
Lastly, let $M_{\alpha,d}=\{\omega^{\alpha+1}\cdot k:1\leq k\leq d\}\subset\omega^{\alpha+1}\cdot d+1$, i.e., the set of maximal rank elements. Then $M_{\alpha,d}$ is preserved by every element of $\operatorname{H}_{\alpha,d}$, yielding an action of this group on $M_{\alpha,d}$—which we call the canonical action. We let $\operatorname{PH}_{\alpha,d}$ denote the kernel of this action, i.e., the subgroup of $\operatorname{H}_{\alpha,d}$ fixing each element of $M_{\alpha,d}$; we call this group the pure homeomorphism group of $\omega^{\alpha}\cdot d+1$. Note that $\operatorname{H}_{\alpha,1}=\operatorname{PH}_{\alpha,1}=\overline{\operatorname{F}}_{\alpha,1}$, where the last equality follows from 3.12.
Recall that a short exact sequence $1\to A\to B\to C\to 1$ is right split if the map $B\to C$ admits a section.
4.4.
@@
[…]
    is a short exact sequence and it is right split.
@@
    If $\Phi\colon\thinspace\operatorname{H}_{\alpha,d}\to\operatorname{Sym}(M_{\alpha,d})$ is the canonical action, then the short exact sequence
    $1\longrightarrow\operatorname{PH}_{\alpha,d}\lhook\joinrel\longrightarrow\operatorname{H}_{\alpha,d}\overset{\Phi}{\longrightarrow}\operatorname{Sym}(M_{\alpha,d})\longrightarrow 1$
    is right split.
A short exact sequence is right split if and only if the middle group is a semi-direct product of the other two groups, yielding the following corollary.
@@
(Semi-direct product decomposition) Let $\alpha$ be an ordinal, and let $d\in\mathbb{N}$. If $d>1$, then $\operatorname{PH}_{\alpha,d}$ is isomorphic to $\overline{\operatorname{F}}_{\alpha,d}\rtimes\mathbb{Z}^{d-1}$ and $\operatorname{H}_{\alpha,d}$ is isomorphic to $\operatorname{PH}_{\alpha,d}\rtimes\operatorname{Sym}(d)$. In particular, $\operatorname{H}_{\alpha,d}$ is isomorphic to $(\overline{\operatorname{F}}_{\alpha,d}\rtimes\mathbb{Z}^{d-1})\rtimes\operatorname{Sym}(d)$.
An explicit description of $\chi$, and its section, is given in Section 4.1, which comes from $d-1$ independent flux homomorphisms counting the shifting of the elements of $N_{\alpha,d}$. As a consequence, we see that $\operatorname{PH}_{\alpha,d}$ and $\operatorname{H}_{\alpha,d}$ are neither perfect nor strongly bounded. For the latter, note that $\operatorname{PH}_{\alpha,d}$ has a countably infinite quotient and hence is not strongly bounded per the discussion above, and as $\operatorname{H}_{\alpha,d}$ has a finite-index subgroup that fails to be strongly bounded, it must also fail to be strongly bounded.
A topological group is coarsely bounded if every continuous left-invariant metric on the group has bounded diameter (see [26]). Equipped with the compact-open topology, $\operatorname{H}_{\alpha,d}$ is a topological group (see Section 2). With this topology, it is readily verified that the homomorphism $\chi$ is continuous, and hence it follows from 4.4 that $\operatorname{PH}_{\alpha,d}$ (and hence $\operatorname{H}_{\alpha,d}$) is not coarsely bounded. Together with 3.14, this yields the following corollary.
@@
Recently, Branman–Domat–Hoganson–Lyman [8, Theorem A] showed that for a countable ordinal $\alpha$, the group $\operatorname{Homeo}(\omega^{\alpha}\cdot d+1)$ is coarsely bounded if and only if $d=1$. 1.3 is a strengthening of this statement under the additional hypothesis that $\alpha$ is a successor ordinal. We do not know if $\operatorname{Homeo}(\omega^{\lambda})$ is strongly bounded when $\lambda$ is a countable limit ordinal; however, we do know that it is coaresly bounded [23].
Before moving on, we stress that Corollary 1.3 considers $\operatorname{H}_{\alpha,d}$ equipped with the compact-open topology. Given any ordinal $\alpha$, Gheysens [15, 16] has proven that $\operatorname{Homeo}(\alpha)$, equipped with the topology of pointwise convergence, is a Roelcke precompact topological group, implying that it is coarsely bounded.
Ideally, we would like to use 4.4 to compute the abelianizations of $\operatorname{PH}_{\alpha,d}$ and $\operatorname{H}_{\alpha,d}$, as well as to produce normal generating sets. The main obstacle is working with $\operatorname{F}_{\alpha,d}$, which is a complicated group. If we quotient out by $\operatorname{F}_{\alpha,d}$, then we do have the tools to proceed. The downside is that the isomorphism class of this quotient does not depend on $\alpha$. Indeed, let $\mathrm{K}_{\alpha,d}$ denote the kernel of the action of $\operatorname{H}_{\alpha,d}$ on the set $N_{\alpha,d}$. It can be readily verified that $\operatorname{H}_{\alpha,d}/\mathrm{K}_{\alpha,d}$ is isomorphic to $\operatorname{H}_{0,d}$, leading us to the investigation of the groups $\operatorname{Homeo}(\omega\cdot d+1)$. The difference in the case $\alpha=0$ is that $\operatorname{F}_{0,d}$ is the group of finitely supported homeomorphisms; in particular, the elements of $\operatorname{F}_{0,d}$ are supported away from the set $M_{0,d}$.
@@
4.9 stated below contains another condition—related to 4.8—that we omit here, as we do not have the appropriate definitions yet at hand. The proof of 4.9 follows from a fragmentation lemma, 4.5, that holds for all $\alpha$ but that is strongest when $\alpha=1$, again as $\operatorname{F}_{0,d}$ consists only of finitely supported homeomorphisms. Fragmentation will allow us to view any element of $\overline{\operatorname{F}}_{0,d}$ as a composition of $d+1$ homeomorphisms, each of which can be viewed as homeomorphism of $\operatorname{Homeo}(\omega)$. This fragmentation also allows to consider the abelianization of $\overline{\operatorname{F}}_{0,d}$
@@
The computation of the abelianization of $\operatorname{PH}_{0,d}$ and of $\operatorname{H}_{0,d}$ implies that the minimal cardinality for a normal generating set is at least $d-1$ and at least two, respectively. Our final theorem, says this is in fact an equality, and the proofs construct explicit normal generating sets of the desired cardinalities.
@@
If $d\in\mathbb{N}\smallsetminus\{1\}$, then the minimal cardinality of a normal generating set for $\operatorname{PH}_{0,d}$ is $d-1$ and is two for $\operatorname{H}_{0,d}$.
@@
Motivation and context
@@
Now, if $E$ is a second-countable Stone space (i.e., $E$ is second countable, compact, zero-dimensional, and Hausdorff), then $E$ embeds as a subset of $\mathbb{S}^{2}$. Identifying $E$ with such an embedding, it follows from [25] that the canonical homomorphism $\operatorname{Homeo}(\mathbb{S}^{2}\smallsetminus E)\to\operatorname{Homeo}(E)$ is surjective. Moreover, as any two isotopic homeomorphisms induce the same action on $E$, this homomorphism factors to give an epimorphism $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)\to\operatorname{Homeo}(E)$, where $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is the mapping class group of $\mathbb{S}^{2}\smallsetminus E$, i.e., the group of isotopy classes of homeomorphisms $\mathbb{S}^{2}\smallsetminus E\to\mathbb{S}^{2}\smallsetminus E$. The group $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is an example of a big mapping class group, a class of groups that have been investigated intensely recently. We therefore see that theorems about big mapping class groups have ramifications for theorems regarding homeomorphism groups of second-countable Stone spaces.
@@
Homeomorphism groups of Stone spaces have a long history of being studied from the perspective of automorphism groups of Boolean algebras, and the mapping class group perspective has brought a renewed interest. Ordinals are Stone spaces, and the topological structure of ordinals are simpler than Stone spaces in general. This has led to a recent interest in homeomorphism groups of ordinals, e.g., see [17, 7, 8].
Many of the theorems here are motivated by work of the last author and his collaborators on big mapping class groups, see [29, 2]. But with that said, we stress that our results do not follow from that setting. For one, only countable successor ordinals can be realized as the end space of a manifold (assuming the manifold is second countable; even dropping the second countability assumption on the manifold, the type of ordinals that can appear are quite limited, see [13] for a discussion of ends in non-metrizable manifolds).

### v2

@@ 1 Introduction
Recall that an ordinal is the order-isomorphism class of a totally ordered set that is well-ordered in the sense that every nonempty subset has a smallest element. Every totally ordered set admits a topology—the order topology—with a subbasis given by the open rays $\{x:x<y\}$ and $\{x:y<x\}$ ranging over the elements $y$ in the set. As an order isomorphism between totally ordered sets induces a homeomorphism between their respective order topologies, an ordinal has a well-defined homeomorphism type, allowing us to study the homeomorphism group of an ordinal.
Motivated by recent results regarding homeomorphism groups of surfaces, we investigate several algebraic and geometric properties of homeomorphism groups of ordinals. We provide a discussion of this motivation from geometric topology after introducing our main results.
For an introduction to ordinals and the terminology that follows, see Section 2. Our results focus on successor ordinals, or equivalently, the compact ordinals. The homeomorphism type of an infinite successor ordinal is classified by two parameters: an ordinal $\alpha$ (called the limit capacity) and a natural number $d$ (called the coefficient). The limit capacity and the coefficient are invariants of a well-ordered set that can be computed topologically as follows: the limit capacity is equal to the predecessor of the Cantor–Bendixson rank and the coefficient is equal to the Cantor–Bendixson degree. We study the successor ordinals whose limit capacity is itself a successor ordinal.
Let us fix some notation. Given an ordinal $\alpha$ and a natural number $d$, we let $X_{\alpha,d}$ denote a well-ordered set representing a compact ordinal whose limit capacity is $\alpha+1$ and whose coefficient is $d$. We now let $\operatorname{H}_{\alpha,d}$ denote $\operatorname{Homeo}(X_{\alpha,d})$, the group consisting of homeomorphisms $X_{\alpha,d}\to X_{\alpha,d}$. In Section 2, we will introduce a canonical choice for $X_{\alpha,d}$ using von Neumann ordinals and the classification of ordinals up to homeomorphism; in particular, in later sections, we set $X_{\alpha,d}=\omega^{\alpha+1}\cdot d+1$.
Our first slate of results is for the case when the coefficient is one. Groups in this class generalize $\operatorname{Sym}(\mathbb{N})$, the symmetric group on the natural numbers, as $\operatorname{Sym}(\mathbb{N})$ is isomorphic to $\operatorname{H}_{0,1}$.
An element $g$ in a group $G$ is a normal generator of $G$ if every element of $G$ can be expressed as a product of conjugates of $g$ and $g^{-1}$; it is a uniform normal generator of $G$ if there exists $k\in\mathbb{N}$ [^1] Throughout, we use the convention that $0\notin\mathbb{N}$. such that every element of $G$ can be expressed as a product of at most $k$ conjugates of $g$ and $g^{-1}$, and the minimum such $k$ is called the $g$-width of $G$.
Given $x\in X_{\alpha,d}$, the set $\{y:y\leq x\}$ is a well-ordered set, allowing us to define the capacity of $x$ to be equal to the limit capacity of $\{y:y\leq x\}$.
@@
Let $\alpha$ be an ordinal. For $h\in\operatorname{H}_{\alpha,1}$, the following are equivalent:
@@
    $h$ normally generates $\operatorname{H}_{\alpha,1}$.
@@
    $h$ induces an infinite permutation of the capacity $\alpha$ elements of $X_{\alpha,1}$.
Moreover, if one—and hence all—of the above conditions are satisfied, then the $h$-width of $\operatorname{H}_{\alpha,d}$ is at most twelve.
The proof of Theorem 3.12 relies on a uniform fragmentation result—a generalization of a lemma of Galvin [12] from permutation groups—and a technique from Anderson’s proof [1] of the algebraic simplicity of several transformation groups, including the homeomorphism group of the Cantor set and the homeomorphism group of the 2-sphere. As an immediate consequence, $\operatorname{H}_{\alpha,1}$ has a maximal proper normal subgroup.
@@
Let $\alpha$ be an ordinal. The subgroup of $\operatorname{H}_{\alpha,1}$ consisting of homeomorphisms that induce finite permutations on the capacity $\alpha$ elements in $X_{\alpha,1}$ is the maximal proper normal subgroup of $H_{\alpha,1}$, that is, it contains every proper normal subgroup. ∎
As $\operatorname{H}_{0,1}$ is isomorphic to $\operatorname{Sym}(\mathbb{N})$, 1.1 recovers the classical theorem of Schreier–Ulam [26] that the subgroup of finite permutations in $\operatorname{Sym}(\mathbb{N})$ is its maximal normal subgroup. The proof presented here is arguably simpler than the standard proof given in classical texts such as [27], which relies on cycle decompositions. Similarly, 3.12 can be viewed as a generalization of a theorem of Bertram [5] showing that all infinite permutations in $\operatorname{Sym}(\mathbb{N})$ are uniform normal generators. Unlike Bertram’s proof, our proof does not rely on cycle decompositions. Bertram proves that any permutation in $\operatorname{Sym}(\mathbb{N})$ can be expressed as a product of four conjugates of any infinite permutation, raising the question of what the optimal constant in 3.12 is.
Recall that the commutator subgroup of a group $G$, denoted $[G,G]$, is the subgroup generated by the set $\{[x,y]:x,y\in G\}$, where $x,y]=xyx^{-1}y^{-1}$. A group $G$ is perfect if it is equal to its commutator subgroup. It is uniformly perfect if there exists $k\in\mathbb{N}$ such that every element of $G$ can be expressed as a product of $k$ commutators; the minimum such $k$ is called the commutator width of $G$.
It is not difficult to find a commutator in $\operatorname{H}_{\alpha,1}$ that induces an infinite permutation of the capacity $\alpha$ elements in $X_{\alpha,1}$; 3.12 then implies that $H_{\alpha,1}$ is a uniformly perfect group of commutator width at most twelve. A direct proof, which is simpler than the proof of 3.12, can be given of this fact that provides a better bound on the commutator width of $G$. This is our second main result.
@@
If $\alpha$ is an ordinal, then $\operatorname{H}_{\alpha,1}$ is uniformly perfect, and the commutator width is at most three.
A theorem of Ore [23] says that every element of $\operatorname{Sym}(\mathbb{N})$ can be expressed as a single commutator. It is natural to ask whether the same is true for $\operatorname{H}_{\alpha,1}$.
Continuing in analogy with $\operatorname{Sym}(\mathbb{N})$, we turn to a geometric property. In [4], Bergman showed that any left-invariant metric on $\operatorname{Sym}(\mathbb{N})$ has bounded diameter; in the literature, this property is referred to as strong boundedness or the Bergman property. There are several equivalent definitions of strongly bounded, one of relevance here is that every group action on a metric space has bounded orbits. As a consequence, any homomorphism from a strongly bounded group to a countable group has finite image, a fact we use below. Note that strong boundedness is a property of a group as a discrete group, so there are no continuity requirements for the metrics, actions, or homomorphisms in the above discussion.
Adapting a technique appearing several places in the literature (see [19, Construction 2.3]), we establish $\operatorname{H}_{\alpha,1}$ is strongly distorted. The notion of strong distortion was introduced by Calegari–Freedman in [8], where they established strong distortion for homeomorphism groups of spheres; in the appendix of the same paper, Cornulier showed that if $G$ is strongly distorted, then $G$ is strongly bounded. See Section 3.4 for the definition of strong distortion.
@@
For every ordinal $\alpha$, the group $\operatorname{H}_{\alpha,1}$ is strongly distorted.
@@
We finish our discussion of the degree one case by observing that we cannot drop the hypothesis that the limit capacity is a successor. It is a consequence of [11, Theorem 3.9] that every countable group is a quotient of $\operatorname{Homeo}(\omega_{1})$, where $\omega_{1}$ is the first uncountable ordinal. As a consequence, $\operatorname{Homeo}(\omega_{1})$ can be neither uniformly perfect, strongly bounded, nor strongly distorted.
We now turn to the case of coefficient, or equivalently Cantor–Bendixson degree, greater than one, investigating the algebraic structure of $\operatorname{H}_{\alpha,d}$ and establishing the failure of the above results when $d>1$. There are several normal subgroups of $\operatorname{H}_{\alpha,d}$ that are essential to our results, which we must introduce.
The first is $\operatorname{F}_{\alpha,d}$, the subgroup consisting of the homeomorphisms of $X_{\alpha,d}$ inducing a finite permutation of the capacity $\alpha$ elements of $X_{\alpha,d}$. The second is obtained by equipping $\operatorname{H}_{\alpha,d}$ with the compact-open topology and taking the closure of $\operatorname{F}_{\alpha,d}$, which we denote $\overline{\operatorname{F}}_{\alpha,d}$. Lastly, we let $\operatorname{PH}_{\alpha,d}$ denote the subgroup of $\operatorname{H}_{\alpha,d}$ fixing each element $X_{\alpha,d}$ of capacity $\alpha+1$. Note that, as there are only $d$ capacity $\alpha+1$ elements, $\operatorname{PH}_{\alpha,d}$ is a finite-index subgroup, and the corresponding quotient group is isomorphic to $\mathrm{Sym}(d)$, the symmetric group on $d$ letters. Observe that $\operatorname{H}_{\alpha,1}=\operatorname{PH}_{\alpha,1}=\overline{\operatorname{F}}_{\alpha,1}$, where the last equality follows from 3.12.
Recall that a short exact sequence $1\to A\to B\to C\to 1$ is split if the map $B\to C$ admits a section. Also note that, equipped with the compact-open topology, $\operatorname{H}_{\alpha,d}$ is a topological group (see Section 2).
$4.$
@@
    $1\longrightarrow\overline{\operatorname{F}}_{\alpha,d}\lhook\joinrel\longrightarrow\operatorname{PH}_{\alpha,d}\overset{\chi}{\longrightarrow}\mathbb{Z}^{d-1}\longrightarrow 1$
    is a split short exact sequence.
@@
    The short exact sequence
    $1\longrightarrow\operatorname{PH}_{\alpha,d}\lhook\joinrel\longrightarrow\operatorname{H}_{\alpha,d}\longrightarrow\operatorname{Sym}(d)\longrightarrow 1$
    is split.
Moreover, all the homomorphisms in the above exact sequences are continuous when $\operatorname{H}_{\alpha,d}$, $\overline{\operatorname{F}}_{\alpha,d}$, and $\operatorname{PH}_{\alpha,d}$ are equipped with the compact-open topology and $\mathbb{Z}^{d-1}$ and $\operatorname{Sym}(d)$ are equipped with the discrete topology.
If a topological group $G$ is abstractly the semidirect product of its subgroups $N$ and $K$, then this semidirect product is topological if both $N$ and $H$ are closed and the map $N\times H\to G$ given by $(n,h)\mapsto nh$ is a homeomorphism.
As $\chi$ is continuous with discrete image, the restriction of $\chi$ to the image of a section is a homeomorphism, implying the image of any section is discrete. Moreover, the continuity of $\chi$ and of the homomorphism $\operatorname{PH}_{\alpha,d}\to\operatorname{Sym}(d)$ implies that $\overline{\operatorname{F}}_{\alpha,d}$ and $\operatorname{PH}_{\alpha,d}$ are clopen subgroups, respectively, from which it readily follows that $\operatorname{PH}_{\alpha,d}$ is homeomorphic to $\operatorname{F}_{\alpha,d}\times\mathbb{Z}^{d-1}$ and $\operatorname{H}_{\alpha,d}$ is homeomorphic to $\operatorname{PH}_{\alpha,d}\times\operatorname{Sym}(d)$.
Recall that a short exact sequence is split if and only if the middle group is a semidirect product of the other two groups. Therefore, 4.4 yields a topological semi-direct product structure for $\operatorname{PH}_{\alpha,d}$ and $\operatorname{H}_{\alpha,d}$, which we record in the following lemma.
@@
(Topological semidirect product decomposition) Let $\alpha$ be an ordinal, and let $d\in\mathbb{N}$ with $d>1$. Both $\operatorname{PH}_{\alpha,d}$ and $\operatorname{H}_{\alpha,d}$, equipped with the compact-open topology, admit topological semidirect product decompositions. Specifically,
$\operatorname{PH}_{\alpha,d}=\overline{\operatorname{F}}_{\alpha,d}\rtimes\mathbb{Z}^{d-1}$
and
$\operatorname{H}_{\alpha,d}=PH_{\alpha,d}\rtimes\operatorname{Sym}(d)=\left(\overline{\operatorname{F}}_{\alpha,d}\rtimes\mathbb{Z}^{d-1}\right)\rtimes\operatorname{Sym}(d),$
where $\mathbb{Z}^{d-1}$ and $\operatorname{Sym}(d)$ are discrete and are identified with their images under the sections from 4.4.
An explicit description of $\chi$, and its section, is given in Section 4.1, which comes from d1d-1 independent flux homomorphisms counting the shifting of the capacity $\alpha$ elements of $X_{\alpha,d}$. As a consequence, we see that $\operatorname{PH}_{\alpha,d}$ and $\operatorname{H}_{\alpha,d}$ are neither perfect nor strongly bounded. For the latter, note that $\operatorname{PH}_{\alpha,d}$ has a countably infinite quotient and hence is not strongly bounded per the discussion above, and as $\operatorname{H}_{\alpha,d}$ has a finite-index subgroup that fails to be strongly bounded, it must also fail to be strongly bounded.
A topological group is coarsely bounded if every continuous left-invariant metric on the group has bounded diameter (see [25]). Note that this is a weaker condition than being strongly bounded, as we have required the metrics under consideration to be continuous; in particular, a group is strongly bounded if and only if it is coarsely bounded as a discrete group. As mentioned in 4.4, the homomorphism $\chi$ is continuous, and hence it follows from 4.4 that $\operatorname{PH}_{\alpha,d}$ (and hence $\operatorname{H}_{\alpha,d}$) is not coarsely bounded. Together with 3.14, this yields the following corollary.
@@
Recently, Branman–Domat–Hoganson–Lyman [7, Theorem A] showed that the homeomorphism group of a countable successor ordinal is coarsely bounded if and only if its coefficient is one. 1.3 is a strengthening of this statement under the additional hypothesis that the limit capacity is a successor ordinal, and it also broadens the statement beyond the countable ordinals. In the case of a countable coefficient one successor ordinal whose limit capacity is a limit ordinal, we do not know if its homeomorphism group is strongly bounded, despite knowing that it is coarsely bounded.
Before moving on, we stress that 1.3 considers $\operatorname{H}_{\alpha,d}$ equipped with the compact-open topology. Given any ordinal $\alpha$, Gheysens [13, 14] has proven that $\operatorname{Homeo}(\alpha)$, equipped with the topology of pointwise convergence, is a Roelcke precompact topological group, implying that it is coarsely bounded.
Ideally, we would like to use 4.4 to compute the abelianizations of $\operatorname{PH}_{\alpha,d}$ and $\operatorname{H}_{\alpha,d}$, as well as to produce normal generating sets. The main obstacle is working with $\operatorname{F}_{\alpha,d}$, which is a complicated group. If we quotient out by $\operatorname{F}_{\alpha,d}$, then we do have the tools to proceed. The downside is that the isomorphism class of this quotient does not depend on $\alpha$. Indeed, let $\mathrm{K}_{\alpha,d}$ denote the kernel of the action of $\operatorname{H}_{\alpha,d}$ on the set of capacity $\alpha$ elements of $X_{\alpha,d}$. It can be readily verified that $\operatorname{H}_{\alpha,d}/\mathrm{K}_{\alpha,d}$ is isomorphic to $\operatorname{H}_{0,d}$, leading us to the investigation of $\operatorname{H}_{0,d}$. The difference in the case $\alpha=0$ is that $\operatorname{F}_{0,d}$ is the group of finitely supported homeomorphisms, as $X_{0,d}$ has exactly $d$ accumulation points. In particular, the elements of $\operatorname{F}_{0,d}$ are supported away from elements of maximal capacity.
@@
    $f$ induces an infinite permutation of the isolated points in any open neighborhood of an accumulation point.
The actual 4.9 stated in Section 4.2 contains another equivalent condition—related to 4.8—that we omit here, as we do not have the appropriate definitions yet at hand. The proof of 4.9 follows from a fragmentation lemma, 4.5, that holds for all $\alpha$ but that is strongest when $\alpha=0$, again as $\operatorname{F}_{0,d}$ consists only of finitely supported homeomorphisms. Fragmentation will allow us to view any element of $\overline{\operatorname{F}}_{0,d}$ as a composition of $d+1$ homeomorphisms, each of which can be viewed as homeomorphism of $X_{0,1}$. This fragmentation also allows us to consider the abelianization of $\overline{\operatorname{F}}_{0,d}$.
@@
The computation of the abelianization of $\operatorname{PH}_{0,d}$ and of $\operatorname{H}_{0,d}$ implies that the minimal cardinality for a normal generating set is at least d1d-1 and at least two, respectively. Our final theorem, says this is in fact an equality, and the proofs construct explicit normal generating sets of the desired cardinalities.
@@
If $d\in\mathbb{N}\smallsetminus\{1\}$, then the minimal cardinality of a normal generating set for $\operatorname{PH}_{0,d}$ is d1d-1 and is two for $\operatorname{H}_{0,d}$.
@@
Connections to geometric topology
As mentioned at the beginning of the introduction, the results above were motivated by recent work in the setting of homeomorphism groups of surfaces. We present this connection here, and also contextualize our results in the surface setting.
@@
Recall that a Stone space is a compact, zero-dimensional, Hausdorff topological space. If $E$ is a second-countable Stone space, then $E$ embeds as a subset of $\mathbb{S}^{2}$. Identifying $E$ with such an embedding, it follows from [24] that the canonical homomorphism $\operatorname{Homeo}(\mathbb{S}^{2}\smallsetminus E)\to\operatorname{Homeo}(E)$ is surjective. Moreover, as any two isotopic homeomorphisms induce the same action on $E$, this homomorphism factors to give an epimorphism $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)\to\operatorname{Homeo}(E)$, where $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is the group of isotopy classes of homeomorphisms $\mathbb{S}^{2}\smallsetminus E\to\mathbb{S}^{2}\smallsetminus E$, known as the mapping class group. If $E$ is infinite, the group $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is an example of a big mapping class group, a class of groups that have been investigated intensely recently. We therefore see that theorems about big mapping class groups yield corollaries for homeomorphism groups of second-countable Stone spaces.
@@
Homeomorphism groups of Stone spaces have a long history of being studied from the perspective of automorphism groups of Boolean algebras, which are dual to Stone spaces by Stone duality. The mapping class group perspective has brought a renewed interest. The compact ordinals are Stone spaces, and the topological structure of ordinals are simpler than Stone spaces in general. This has led to a recent interest in homeomorphism groups of ordinals, e.g., see [15, 6, 7].